df to.10 0.025 t0.01 to.0ns df Values of ta 3.078 6.314 2.706 31.821 63.657 1886 2.920 4.303 6.965 9.925 1.638 2.353 3.182 .54 5.841 1.533 2.132 2.776 3.747 4.604 0 1.476 2.015 2.5 3.365 4.032 1.440 1.93 2.4473.143 3.707 1.415 1895 2.365 2.998 3.499 1.397 1860 2.306 2.896 3.355 1.383 1833 2.22 2.82 3.250 1.372 1.812 .282.764 3.169 1.363 1.796 2.20 2.718 3.106 1.356 1.782 2.179 2.68 3.055 1.350 77 2.160 2.650 3.012 1.345 76 2.145 2.624 2.977 10 10 12 13 14 15 16 17 18 19 20 21 12 13 14 15 16 17 18 19 20 21 1.341 1.753 2.3 2.602 2.947 1.337 1746 2.120 2.583 2.921 1.333 1.740 2.110 2.567 2.898 1.330 734 2.10 2.552 2.878 1.328 1.729 2.093 2.539 2.861 1.325 1725 2.086 2.528 2.845 1.323 172 2.080 2.518 2.83 1.32 1717 2.074 2.508 2.819 1.319 71 2.069 2.500 2.807 1.318 7 2.4 2.492 2.797 23 24 23 241-294 1.293 666 993 2.379 2.646 1.293 .666 993 2.379 2.645 1.293 666 993 2.378 2.644 1.293 1665 992 2.377 2.643 1.292 664 990 2.374 2.639 1.292 663 988 2.37 2.635 1.29 662 987 2.368 2.632 1.29 66 985 2.366 2.629 1.290 1660 984 2.364 2.626 1.286 653 972 2.345 2.601 1.284 1.650 968 2.339 2.592 1.284 649 966 2.336 2.588 1.283 1.648 965 2.334 2.586 1.283 647 964 2.333 2.584 1.283 647 1.963 2.332 2.583 1.283 647 963 2.33 2.582 1.282 647 963 2.330 2.581 1.282 1646 1962 2.330 2.581 1.282 1646 196 2.328 2.578 1.994 264 72 73 74 72 3 74 75 80 95 100 200 300 400 500 95 100 200 300 400 500 700 700 800 1000 2000 1000 2000 df 0.05 o.02s o.0 df о, 10 0,05 0.,005 1.282 1645 960 2.3262.576 2 0.10 0.0s Z 0.025 2 0.01 2 0.005Use the one-mean t-interval procedure with the sample mean, sample size, sample standard deviation, and confidence level given below to find a confidence interval for the mean of the population from which the sample was drawn. X- 2.0 Click here to view page 1 of the table of critical values for the t distribution s 4.7 confidence level-98% Click here to view page 2 of the table The 98% confidence interval about is to (Round to two decimal places as needed.) of tical values fo r the t distribution
Answer #1

Given :-

Sample mean ( large ar X ) = 2.0

Sample standard deviation (s) = 4.7

Sample size (n) = 61

Degree of freedom (df) =n-1 =61-1 =60

Confidence level (c) = 98%

Then critical value is t​​​c= 2.39

Now, we have to find confidence interval for mean

Therefore,

  large mu = ( large ar X pm t_c * (s/sqrt n ) )

large mu = (  2.0± 2.39 * (47N61) )

large mu = ( 2.0 - 1.44 , 2.0 + 1.44 )

  large mu​​​​​​ = ( 0.56 , 3.44 )

Therefore the 98% confidence interval about mean (large mu) is 0.56 to 3.44 .